# The wanted extension of Fujii and Tsurumaru’s formula for the spectral radius of the Bell-CHSH operator

### v1

@@
Abstract. This paper is motivated by a recent paper of Yuki Fujii and Toyohiro Tsurumaru in which they established a beautiful formula for the spectral radius of the Bell-CHSH operator on finite-dimensional Hilbert spaces. To tackle the operator on infinite-dimensional spaces, they elaborated a method based on appropriate approximation of commutators of infinite-dimensional orthogonal projections by commutators of orthogonal projections on finite-dimensional spaces. We here give a proof of Fujii and Tsurumaru’s original formula that works in all dimensions. We also simplify and correct some arguments employed by Fujii and Tsurumaru.
MSC. Primary: 47A25; Secondary: 20C07, 47B35, 47B47, 47B93
Key words. Two orthogonal projections, Commutator, Spectral radius, Bell-CHSH operator
@@
on $V_{1}\otimes V_{2}$. See, e.g., [4, 5, 8, 9]. It is known that the spectral radius of $\mathcal{B}$} can be computed by the formula
@@
where [A,B]=ABBA[A,B]=AB-BA. Since $A_{i},B_{i}]=4[P_{i},Q_{i}]$, we are led to finding an expression for the spectral radius of the commutator of two orthogonal projections.
Thus, let $P,Q$ be two orthogonal projections on a separable complex Hilbert space $V$ and put A=2PI,B=2QIA=2P-I,B=2Q-I.
Theorem 1
We always have $\varrho([P,Q])\leq 1/2$ and $\varrho([A,B])\leq 2$.
Proof. By Halmos’ theorem [7] as quoted in Theorem 1.2 of [2], the commutator $[P,Q]$ is unitarily similar to $(0,0,0,0)\oplus\begin{pmatrix}0&W\\ -W&0\end{pmatrix}$ where $W=\sqrt{H(I-H)}$ with a selfadjoint operator $H$ such that $0\leq H\leq 1$. Taking the determinant of $\begin{pmatrix}-\lambda I&W\\ -W&-\lambda I\end{pmatrix}$, is easily seen that the spectral radius of $[P,Q]$ is equal to the spectral radius of $W$, and the latter is $\sqrt{x(1-x)}$ where $x$ is the point in the spectrum of $H$ that is closest to $1/2$. Consequently, we always have $\varrho([P,Q])\leq 1/2$. Since $[A,B]=4[P,Q]$, we get $\varrho([A,B])\leq 2$.    $\square$
Theorem 2
(Tsirelson’s inequality [10]). We always have $\varrho({\mathcal{B}})\leq 2\sqrt{2}$.
Proof. From Theorem 1 and (1) we infer that $\varrho({\mathcal{B}})\leq\sqrt{4+2\cdot 2}=2\sqrt{2}$.   $\square$
Theorem 3
We have $\varrho([A,B])=\sqrt{\lambda^{2}(4-\lambda^{2})}$ where $\lambda$ in $\sigma(A+B)$, the spectrum of $A+B$, is such that $\lambda^{2}/4$ is closest to $1/2$.
Proof. Let us first assume that $P$ and $Q$ are in general position, which means that the terms $(1,1,0,0)$ and $(1,0,1,0)$ in Halmos’ Theorem 1.2 of [2] are absent. Then $A+B-\lambda I$ is unitarily similar to $\begin{pmatrix}2(I-H)-\lambda I&2W\\ 2W&2(H-I)-\lambda I\end{pmatrix}$. By computing the determinant of this matrix we see that $\lambda$ is in the spectrum of $A+B$ if and only if $1-\lambda^{2}/4$ is in the spectrum of $H$. The spectrum of $W^{2}=H(I-H)$ is therefore the set of
@@
where $\lambda$ ranges over the spectrum of $A+B$. The spectrum of $[P,Q]$ is the set of $\nu$ for which $\det\begin{pmatrix}-\nu I&W\\ -W&-\nu I\end{pmatrix}=\nu^{2}I+W^{2}$ is not invertible. These values $\nu$ are purely imaginary and we may write $\nu=i\mu$ with real $\mu$. Consequently, the spectrum of $[P,Q]$ is the set of $\nu$ such that $\nu^{2}=\mu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $A+B$. The maximum of $|\mu|$ is attained at a $\lambda$ for which $\lambda^{2}/4$ is closest to $1/2$. For this $\lambda$, we have
$\varrho([A,B])=4\varrho([P,Q])=4|\mu|=4\sqrt{\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)}=\sqrt{\lambda^{2}(4-\lambda^{2})},$
@@
If $P,Q$ are not in general position, then $[P,Q]$ is as in the proof of Theorem 1 and hence the spectrum of $[P,Q]$ is the union of $\{0\}$ and the above set of $\mu$ for which $\mu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $\begin{pmatrix}2(I-H)&2W\\ 2W&2(H-I)\end{pmatrix}$, and we therefore get the same expression for $\varrho([A,B])$.    $\square$
For finite-dimensional spaces $V$, this theorem was established by Fujii and Tsurumaru in [6]. They wrote “when $V$ is infinite-dimensional, it is not known whether or how the equality $\varrho([A,B])=\sqrt{\lambda^{2}(4-\lambda^{2})}$ can be extended, either in the form of an equality or an inequality.” Theorem 3 is the extension they were asking about.
@@
In [6] it is claimed that the $\lambda$ in the formula $\varrho([A,B])=\sqrt{\lambda^{2}(4-\lambda^{2})}$ is the $\lambda$ closest to $\sqrt{2}$. However, there is a subtlety in the matter: a $\lambda$ closest to $\sqrt{2}$ need not be a $\lambda$ for which $\lambda^{2}/4$ is closest to $1/2$. Indeed, if, for example, $A$ and $B$ are given on ${\mathbf{C}}^{3}$ by the involutive matrices
@@
with $x=0.01$, then the spectrum $\sigma(A+B)$ equals $\{-2\sqrt{x},2\sqrt{x},2\}=\{-0.2,0.2,2\}$, the point $\lambda=2$ is closest to $\sqrt{2}$, but $\lambda=\pm 0.2$ are the points for which $\lambda^{2}/4$ is closest to $1/2$.
Let $\theta\in(0,\pi)$ and put $s=\sin\theta$ and $c=\cos\theta$. Note that $c^{2}+s^{2}=1$ and thus $\frac{1-c}{s}=\frac{s}{1+c}$ and $\frac{1+c}{s}=\frac{s}{1-c}$. Furthermore, if $\theta<\pi/2$ then $c+s>1$, and if $\theta>\pi/2$ then 1<c+s<1-1<c+s<1. Fujii and Tsurumaru consider a pair of involutions in so-called constant-angle one-shifted form. Such a pair is given by
@@
The following theorem is already in [6], but the proof given there is based on the incorrect assumption that $\sigma(A(\theta)+B(\theta))=[-2s,2s]$ for $\theta\in(0,\pi/4)$, which mistake was fortunately neutralized by the fact that the outlier $\lambda=2$, though being closest to $\sqrt{2}$, is not the $\lambda$ for which $\lambda^{2}/4$ is closest to $1/2$.
Theorem 5
@@
Proof. We use Theorems 3 and 4. If $\theta\leq\pi/4$, then $\lambda^{2}/4=2^{2}/4=1$ has distance $1/2$ to $1/2$ whereas $\lambda^{2}/4=(2s)^{2}/4=s^{2}$ is at the distance $1/2-s^{2}$ to $1/2$. Consequently, $\lambda=2s$ is a point in the spectrum for which $\lambda^{2}/4$ is closest to $1/2$. This implies that
@@
If $\pi/4\leq\theta\leq 3\pi/4$, then $\sqrt{2}\in[-2s,2s]$ and hence $\lambda=\sqrt{2}$ minimizes the distance between $\lambda^{2}/4$ and $1/2$. Thus, $\varrho([A(\theta),B(\theta)])=\sqrt{2(4-2)}=2$. Finally, if $3\pi/4\leq\theta<\pi$, then again $\lambda^{2}/4=(2s)^{2}/4=s^{2}$ is closest to $1/2$, and since $c<0$ in this case, we obtain $\varrho([A(\theta),B(\theta)])=2|\sin(2\theta)|$.    $\square$
@@
where $r(\eta)=\begin{pmatrix}\cos\eta&\sin\eta\\ \sin\eta&-\cos\eta\end{pmatrix}$ and $\{\omega_{j}\}_{=1}^{\infty},\{\theta_{j}\}_{=1}^{\infty}$ are sequences of points in $(0,\pi)$. We say that two such involutions are in eventually constant-angle one-shifted form if there is a $j_{0}$ such that $\omega_{j}=\theta_{j}=\theta$ for all $j\geq j_{0}$. Let $P_{n}$ be the projection that sends a sequence $\{x_{j}\}_{j=0}^{\infty}$ to the sequence $\{x_{0},\ldots,x_{n-1},0,0,\ldots\}$. If $T$ is an operator on $\ell^{2}({\mathbf{Z}}_{+})$, then $P_{n}TP_{n}$ may be identified with an $n\times n$ matrix in the natural fashion. The spectrum $\sigma(P_{n}TP_{n})$ is nothing but the set of the eigenvalues of the matrix $P_{n}TP_{n}$.
Theorem 6
If $A$ and $B$ are in eventually constant-angle one-shifted form, then the spectrum of $P_{n}(A+B)P_{n}$ converges in the Hausdorff metric to the spectrum of $A+B$.
Proof. In the case at hand, we have $A+B=T(a)+K$ where, as above, $T(a)$ is the infinite tridiagonal Toeplitz matrix whose symbol is $a(e^{i\varphi})=2s\cos\varphi$ with $s=\sin\theta$ and $K$ is a finite-rank operator with a finite matrix in the upper left corner and zeros elsewhere. By Proposition 2.33 and Theorem  4.16 of [1], the spectrum of $P_{n}(T(a)+K)P_{n}$ converges in the Hausdorff metric to $\sigma(T(a)+K)\cup\sigma(T(a))$. (Note that $T(a)$ is symmetric, which implies that the T(a~)T($\widetilde${a}) occurring in Theorem 4.16 coincides with $T(a)$.) Since $\sigma(T(a))=[-2s,2s]$ and $\sigma(T(a)+K)$ contains the essential spectrum $\sigma_{\rm ess}(T(a))=[-2s,2s]$, we arrive at the assertion.    $\square$
@@
Figure 1: The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=10$ (left) and $n=600$ (right) with $A,B$ as in (3) and with $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$.
Instead of solving this equation, let us use Theorem 6. Figure 1 shows the result for $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$. To be sure that the outliers are not produced by numerical instabilities, let us check whether (4) is true for the rational numbers c=0.8c=-0.8 and $s=0.6$, which corresponds to $\theta=\arccos(-0.8)\approx 2.4981$ and where exact computations can be done by hand. The picture suggests that in this case the two outliers are $\lambda=\pm 1.5$. The roots of $q^{2}-(1.5/0.6)q+1$ are $2$ and $1/2$. Hence, we have to take $q=0.5$. We have $\gamma=0$ and $\delta=1$ for $\omega=\pi/2$. Inserting these values in the left-hand side of (4), we obtain indeed 0. The same happens for $\lambda=-1.5$. Thus, we may conclude that $\sigma(A+B)=[-1.2,1.2]\cup\{-1.5,1.5\}$ for $\omega=\pi/2$ and $\theta\approx 2.4981$. As $1.5^{2}/4-0.5=0.0625<0.14=0.5-0.6^{2}$, we see that in the case at hand the spectral radius $\varrho([A,B])$ is not determined by $s$ but rather by the outliers $\lambda=\pm 1.5$. Theorem 3 gives $\varrho([A,B])=\sqrt{1.5^{2}(4-1.5^{2})}\approx 1.9843$. The value of $\sqrt{4s^{2}(4-4s^{2})}$ is $1.9200$.
Figure 2: Plots of $\lambda_{\max}(\theta)$ where $\lambda_{\max}(\theta)$ is the largest eigenvalue of $P_{n}(A+B)P_{n}$ for $n=100$ with $A,B$ as in (3). In the left picture we took $\omega=\pi/2$, in the right picture we have $\omega=0.1:0.1:3.1$. The red curve is $2\sin\theta$, the green curve is $2|\cos\theta|$. The horizontal lines are at the height $\sqrt{2}$. The rightmost vertical line in the left picture has the abscissa $x_{3}=2.4352$.
The blue curve in the left picture of Figure 2 shows the outliers $\lambda>0$ for $\omega=\pi/2$ and $0<\theta<\pi$. The curve is of constant height $2$ for $0<\theta\leq\pi/2$ and then goes down for $\pi/2<\theta<\pi$. The abscissas of the intersections of the red curve $\theta\mapsto 2\sin\theta$ and the green curve $\theta\mapsto|2\cos\theta|$ are $x_{1}=\pi/4$ and $x_{2}=3\pi/4\approx 2.3562$, and the abscissa of the intersection of the green curve and the blue outlier curve $\theta\mapsto\lambda_{\max}(\theta)$ is $x_{3}\approx 2.4352$. For $\theta>x_{2}$, we have $\lambda>\sqrt{2}$ and $2s<\sqrt{2}$, i.e., both $\lambda^{2}/4-1/2$ and $1/2-s^{2}$ are positive. It follows that $\lambda^{2}/4-1/2<1/2-s^{2}$ if and only if $\lambda^{2}<4(1-s^{2})=4c^{2}$. The latter is the case exactly if the blue curve is below the green curve, that is, for $\theta>x_{3}$. In summary, taking for each $\theta$ the value of $\lambda\in\sigma(A+B)$ for which $\lambda^{2}/4$ is closest to $1/2$, we arrive at the following:

### v3

@@
Abstract.
This paper is motivated by a recent paper of Yuki Fujii and Toyohiro Tsurumaru in which they established a beautiful formula for the spectral radius of the Bell-CHSH operator on finite-dimensional Hilbert spaces. To tackle the operator on infinite-dimensional spaces, they elaborated a method based on appropriate approximation of commutators of infinite-dimensional orthogonal projections by commutators of orthogonal projections on finite-dimensional spaces. We here give a proof of Fujii and Tsurumaru’s original formula that works in all dimensions. We also present an alternative approximation procedure, uncover the connection of the problem with block Toeplitz operators, and derive good estimates and explicit expressions for the spectral radius in concrete cases.
Key words and phrases:
Two orthogonal projections, Commutator, Spectral radius, Bell-CHSH operator
1991 Mathematics Subject Classification:
Primary 47A25; Secondary 20C07, 47B35, 47B47, 47B93
1. The Bell-CHSH operator
@@
on $V_{1}\otimes V_{2}$. The spectral radius $\varrho({\mathcal{B}})$ is of interest in quantum mecanics. See, e.g., [4, 5, 6, 7, 9, 10, 12]. Taking into account that $A_{i}$ and $B_{i}$ are involutions, $A_{i}^{2}=B_{i}^{2}=I$, a simple straightforward computation gives
${\mathcal{B}}^{2}=4(I\otimes I)+[A_{1},B_{1}]\otimes[B_{2},A_{2}],$
where [A,B]=ABBA[A,B]=AB-BA. Consequently, since $\varrho(S)=\varrho(-S)$ for every operator $S$, we get the well-known formula
@@
As $A_{i},B_{i}]=4[P_{i},Q_{i}]$, we are so led to the study of the spectral radius of the commutator of two selfadjoint involutions or equivalently, of two orthogonal projections.
Thus, let $P,Q$ be two orthogonal projections on a separable complex Hilbert space $V$ and put A=2PI,B=2QIA=2P-I,B=2Q-I. We have $\|A\|^{2}=\|AA^{*}\|=\|A^{2}\|=1$ and, analogously, $\|B\|=1$. It follows that
$\varrho([A,B])\leq\|[A,B]\|\leq\|A\|\,\|B\|+\|B\|\,\|A\|=2,$
and hence (1) implies that $\varrho({\mathcal{B}})\leq\sqrt{4+2\cdot 2}=2\sqrt{2}$, which is Tsirelson’s famous inequality [12]. We here are interested in expressions for $\varrho([P,Q])$ and thus for $\varrho([A,B])=4\varrho([P,Q])$ and $\varrho({\mathcal{B}})$ that allow us the exact computation of these quantities for concrete operators $P,Q$.
2. An exact formula for the spectral radius
For a bounded linear operator $T$, we denote by $\sigma(T)$ the spectrum of $T$.
Theorem 1.
We have
$\varrho([A,B])=\max_{\lambda\in\sigma(A+B)}\sqrt{\lambda^{2}(4-\lambda^{2})}=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})},$
where $\lambda_{0}$ in $\sigma(A+B)$ is a point for which $\lambda^{2}$ is closest to $2$.
Proof. We use Halmos’ theorem [8] as quoted in Theorem 1.2 of [2]. Let us first assume that $P$ and $Q$ are in general position, which means that the terms $(1,1,0,0)$ and $(1,0,1,0)$ in Theorem 1.2 of [2] are absent. Then $A+B-\lambda I$ is unitarily similar to $\begin{pmatrix}2(I-H)-\lambda I&2W\\ 2W&2(H-I)-\lambda I\end{pmatrix}$. By computing the determinant of this matrix we see that $\lambda$ is in the spectrum of $A+B$ if and only if $1-\lambda^{2}/4$ is in the spectrum of $H$. The spectrum of $W^{2}=H(I-H)$ is therefore the set of
@@ 2. An exact formula for the spectral radius
where $\lambda$ ranges over the spectrum of $A+B$. The spectrum of $[P,Q]$ is the set of $\nu$ for which $\det\begin{pmatrix}-\nu I&W\\ -W&-\nu I\end{pmatrix}=\nu^{2}I+W^{2}$ is not invertible. These values $\nu$ are purely imaginary and we may write $\nu=i\mu$ with real $\mu$. Consequently, the spectrum of $[P,Q]$ is the set of $\nu$ such that $\nu^{2}=\mu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $A+B$. The maximum of $|\mu|$ is attained at a $\lambda_{0}$ for which $\lambda^{2}/4$ is closest to $1/2$ or equivalently, at a $\lambda_{0}$ for which $\lambda^{2}$ is closest to $2$. For this $\lambda_{0}$, we have
$\varrho([A,B])=4\varrho([P,Q])=4|\mu|=4\sqrt{\frac{\lambda_{0}^{2}}{4}\left(1-\frac{\lambda_{0}^{2}}{4}\right)}=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})},$
@@
If $P,Q$ are not in general position, then $[P,Q]$ is unitarily similar to
$(0,0,0,0)\oplus\begin{pmatrix}0&W\\ -W&0\end{pmatrix}$ (2)
and hence the spectrum of $[P,Q]$ is the union of $\{0\}$ and the above set of $\nu$ for which $\nu^{2}=\frac{\lambda^{2}}{4}\left(1-\frac{\lambda^{2}}{4}\right)$ with $\lambda$ in the spectrum of $\begin{pmatrix}2(I-H)&2W\\ 2W&2(H-I)\end{pmatrix}$, and we therefore get the same expression for $\varrho([A,B])$.    $\square$
For finite-dimensional spaces $V$, this theorem was established by Fujii and Tsurumaru in [6]. They wrote “when $V$ is infinite-dimensional, it is not known whether or how the equality $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$ can be extended, either in the form of an equality or an inequality.” Theorem 1 is the extension they were asking about.
@@
Incidentally, and this is again in the vein of [3], one may also employ Halmos’ theorem in its full generality to prove Tsirelson’s inequality. Indeed, we know that $[P,Q]$ is unitarily similar to the operator (2), it is easily seen that the spectral radius of this operator is equal to the spectral radius of $W$, and the latter is $\sqrt{x(1-x)}$ where $x$ is the point in the spectrum of $H$ that is closest to $1/2$. Consequently, we always have $\varrho([P,Q])\leq 1/2$, whence $\varrho([A,B])\leq 2$ and $\varrho({\mathcal{B}})\leq 2\sqrt{2}$.
In [6] it is claimed that the $\lambda_{0}$ in the formula $\varrho([A,B])=\sqrt{\lambda_{0}^{2}(4-\lambda_{0}^{2})}$ is the $\lambda_{0}$ closest to $\sqrt{2}$. However, there is a subtlety in the matter: a $\lambda_{0}$ closest to $\sqrt{2}$ need not be a $\lambda_{0}$ for which $\lambda^{2}$ is closest to $2$. Indeed, if, for example, $A$ and $B$ are given on ${\mathbf{C}}^{3}$ by the selfadjoint and involutive matrices
@@
with $x=0.01$, then the spectrum $\sigma(A+B)$ equals $\{-2\sqrt{x},2\sqrt{x},2\}=\{-0.2,0.2,2\}$, the point $\lambda=2$ is closest to $\sqrt{2}$, but $\lambda=\pm 0.2$ are the points for which $\lambda^{2}$ is closest to $2$.
3. Pairs in convergent-angle one-shifted form
Let $\theta\in(0,\pi)$ and put $s=\sin\theta$ and $c=\cos\theta$. Note that $c^{2}+s^{2}=1$ and thus $\frac{1-c}{s}=\frac{s}{1+c}$ and $\frac{1+c}{s}=\frac{s}{1-c}$. Furthermore, if $\theta<\pi/2$ then $c+s>1$, and if $\theta>\pi/2$ then 1<c+s<1-1<c+s<1. Fujii and Tsurumaru consider pairs of involutions in so-called constant-angle one-shifted form. Such a pair is given by
@@ 3. Pairs in convergent-angle one-shifted form
The essential spectrum $\sigma_{\rm ess}(T)$ is the set of $\lambda\in{\mathbf{C}}$ for which $T-\lambda I$ is not Fredholm, that is, not invertible modulo compact operators.
@@
The following theorem is already in [6], but the proof given there is based on the incorrect assumption that $\sigma(A(\theta)+B(\theta))=[-2s,2s]$ for $\theta\in(0,\pi/4)$, which mistake was fortunately neutralized by the fact that the outlier $\lambda=2$, though being closest to $\sqrt{2}$, is not the $\lambda$ for which $\lambda^{2}$ is closest to $2$.
Theorem 3.
@@
Proof. We use Theorems 1 and 2. If $\theta\leq\pi/4$, then $\lambda^{2}=2^{2}=4$ has distance $2$ to $2$ whereas $\lambda^{2}=(2s)^{2}=4s^{2}$ is at the distance $2-4s^{2}$ to $2$. Consequently, $\lambda_{0}=2s$ is a point in the spectrum for which $\lambda^{2}$ is closest to $2$. This implies that
@@
If $\pi/4\leq\theta\leq 3\pi/4$, then $\sqrt{2}\in[-2s,2s]$ and hence $\lambda_{0}=\sqrt{2}$ minimizes the distance between $\lambda^{2}$ and $2$. Thus, $\varrho([A(\theta),B(\theta)])=\sqrt{2(4-2)}=2$. Finally, if $3\pi/4\leq\theta<\pi$, then again $\lambda_{0}^{2}=(2s)^{2}=4s^{2}$ is closest to $2$, and since $c<0$ in this case, we obtain $\varrho([A(\theta),B(\theta)])=2|\sin(2\theta)|$.    $\square$
@@
where $r(\eta)=\begin{pmatrix}\cos\eta&\sin\eta\\ \sin\eta&-\cos\eta\end{pmatrix}$ and $\{\omega_{j}\}_{=1}^{\infty},\{\theta_{j}\}_{=1}^{\infty}$ are sequences of points in $(0,\pi)$. We say that two such involutions are in convergent-angle one-shifted form if there is a $\theta\in(0,\pi)$ such that $\omega_{j}\to\theta$ and $\theta_{j}\to\theta$ as $j\to\infty$. Let $P_{n}$ be the projection that sends a sequence $\{x_{j}\}_{j=0}^{\infty}$ to the sequence $\{x_{0},\ldots,x_{n-1},0,0,\ldots\}$. If $T$ is an operator on $\ell^{2}({\mathbf{Z}}_{+})$, then $P_{n}TP_{n}$ may be identified with an $n\times n$ matrix in the natural fashion. The spectrum $\sigma(P_{n}TP_{n})$ is nothing but the set of the eigenvalues of the matrix $P_{n}TP_{n}$.
Theorem 4.
If $A$ and $B$ are in convergent-angle one-shifted form, then the spectrum of $P_{n}(A+B)P_{n}$ converges in the Hausdorff metric to the spectrum of $A+B$.
Proof. In the case at hand, we have $A+B=T(a)+K$ where, as above, $T(a)$ is the infinite tridiagonal Toeplitz matrix whose symbol is $a(e^{i\varphi})=2s\cos\varphi$ with $s=\sin\theta$ and $K$ is a compact operator (given by a tridiagonal matrix whose entries go to zero along each diagonal). By Proposition 2.33 and Theorem  4.16 of [1], the spectrum of $P_{n}(T(a)+K)P_{n}$ converges in the Hausdorff metric to $\sigma(T(a)+K)\cup\sigma(T(a))$. (Note that $T(a)$ is symmetric, which implies that the T(a~)T($\widetilde${a}) occurring in Theorem 4.16 coincides with $T(a)$.) Since $\sigma(T(a))=[-2s,2s]$ and $\sigma(T(a)+K)$ contains the essential spectrum $\sigma_{\rm ess}(T(a))=[-2s,2s]$, we arrive at the assertion.    $\square$
@@
Figure 1. The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=10$ (left) and $n=600$ (right) with $A,B$ as in (3) and with $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$. The solid curves are plots of $\pm 2\sin\theta$.
Figure 2. Plots of $\lambda_{\max}(\theta)$ where $\lambda_{\max}(\theta)$ is the largest eigenvalue of $P_{n}(A+B)P_{n}$ for $n=100$ with $A,B$ as in (3). In the left picture we took $\omega=\pi/2$, in the right picture we have $\omega=0.1:0.1:3.1$. The red curve is $2\sin\theta$, the green curve is $2|\cos\theta|$. The horizontal lines are at the height $\sqrt{2}$. The rightmost vertical line in the left picture has the abscissa $x_{3}=2.4352$.
Instead of solving this equation, let us use Theorem 4. Figure 1 shows the result for $\omega=\pi/2$ and $\theta=0.1:0.1:3.1$. To be sure that the outliers are not produced by numerical instabilities, let us check whether (4) is true for the rational numbers c=0.8c=-0.8 and $s=0.6$, which corresponds to $\theta=\arccos(-0.8)\approx 2.4981$ and where exact computations can be done by hand. The picture suggests that in this case the two outliers are $\lambda=\pm 1.5$. The roots of $q^{2}-(1.5/0.6)q+1$ are $2$ and $1/2$. Hence, we have to take $q=0.5$. We have $\gamma=0$ and $\delta=1$ for $\omega=\pi/2$. Inserting these values in the left-hand side of (4), we obtain indeed 0. The same happens for $\lambda=-1.5$. Thus, we may conclude that $\sigma(A+B)=[-1.2,1.2]\cup\{-1.5,1.5\}$ for $\omega=\pi/2$ and $\theta\approx 2.4981$. As $1.5^{2}-2=0.25<0.56=2-4\cdot 0.6^{2}$, we see that in the case at hand the spectral radius $\varrho([A,B])$ is not determined by $s$ but rather by the outliers $\lambda=\pm 1.5$. Theorem 1 gives $\varrho([A,B])=\sqrt{1.5^{2}(4-1.5^{2})}\approx 1.9843$. The value of $\sqrt{4s^{2}(4-4s^{2})}$ is $1.9200$.
Figure 3. The spectra of $P_{n}(A+B)P_{n}$ multiplied by $i$ for $n=100$ with $A,B$ as in (5) and $\theta=0.1:0.1:3.1$.
The blue curve in the left picture of Figure 2 shows the outliers $\lambda>0$ for $\omega=\pi/2$ and $0<\theta<\pi$. The curve is of constant height $2$ for $0<\theta\leq\pi/2$ and then goes down for $\pi/2<\theta<\pi$. The abscissas of the intersections of the red curve $\theta\mapsto 2\sin\theta$ and the green curve $\theta\mapsto|2\cos\theta|$ are $x_{1}=\pi/4$ and $x_{2}=3\pi/4\approx 2.3562$, and the abscissa of the intersection of the green curve and the blue outlier curve $\theta\mapsto\lambda_{\max}(\theta)$ is $x_{3}\approx 2.4352$. For $\theta>x_{2}$, we have $\lambda>\sqrt{2}$ and $2s<\sqrt{2}$, i.e., both $\lambda^{2}-2$ and $2-4s^{2}$ are positive. It follows that $\lambda^{2}-2<2-4s^{2}$ if and only if $\lambda^{2}<4(1-s^{2})=4c^{2}$. The latter is the case exactly if the blue curve is below the green curve, that is, for $\theta>x_{3}$. In summary, taking for each $\theta$ the value of $\lambda_{0}\in\sigma(A+B)$ for which $\lambda^{2}$ is closest to $2$, we arrive at the following: